OpenStudy (anonymous):

Solve by taking the square roots of both sides. (x + 4)2 - 5 = 2

OpenStudy (anonymous):

Simplifying (x + 4) * 2 + -5 = 2 Reorder the terms: (4 + x) * 2 + -5 = 2 Reorder the terms for easier multiplication: 2(4 + x) + -5 = 2 (4 * 2 + x * 2) + -5 = 2 (8 + 2x) + -5 = 2 Reorder the terms: 8 + -5 + 2x = 2 Combine like terms: 8 + -5 = 3 3 + 2x = 2 Solving 3 + 2x = 2 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + 2x = 2 + -3 Combine like terms: 3 + -3 = 0 0 + 2x = 2 + -3 2x = 2 + -3 Combine like terms: 2 + -3 = -1 2x = -1 Divide each side by '2'. x = -0.5 Simplifying x = -0.5

OpenStudy (anonymous):

Hope this helped

OpenStudy (anonymous):

thanks it did